OpenStudy (anonymous):

a(3tx-2b)=c(dx-2) solve equation for x

OpenStudy (anonymous):

apply distributive property on both side of the equation... then combine like terms (all terms with x)... can you do it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3txa-2ba=dxc-2c i think i did it right

OpenStudy (anonymous):

ok continue...

OpenStudy (anonymous):

i dont know what else to do

OpenStudy (anonymous):

put all terms with x on the left-hand side, others on the right-hand side... you know how to do it? you know how to transpose from one side to the other?

OpenStudy (anonymous):

yeah ill try

OpenStudy (anonymous):

3txa-dxc=2ba-2c i think its right

OpenStudy (anonymous):

ok that's right... this time, factor out x on the left-hand side...

OpenStudy (anonymous):

i dont get what you mean

OpenStudy (anonymous):

meaning you pull-out x in each term on the left-hand side... for example... mx+nx = x (m+n)...

OpenStudy (anonymous):

x2(3ta-dc)=2ba-2c?

OpenStudy (anonymous):

where 2 come from on the left-hand side? maybe that is on the right-hand side because it is common to 2ba and -2c...

OpenStudy (anonymous):

its suppose to be x squared

OpenStudy (anonymous):

you mean the problem goes this way?\[a(3tx^2-2b)=c(dx^2-2)\]

OpenStudy (anonymous):

i guess

OpenStudy (anonymous):

why guess? don't you have the original question?

OpenStudy (anonymous):

yeah but the x are not squared

OpenStudy (anonymous):

for the meantime, forget the x squared... if x is not squared...\[x(3ta-cd)=2(ab-c)\]... to solve for x, we cross-multiply, you know how to do it?

OpenStudy (anonymous):

@ShemKingg?